OpenStudy (anonymous):

In the last 2 weeks, a plant grew 1/8 in. and 3/8 in. respectively. How many inches did the plant grow in the last 2 weeks? Express your answer in lowest terms. um o-o well....im not sure if im doing it right.... plz help :3

OpenStudy (anonymous):

:3

OpenStudy (anonymous):

o-o

OpenStudy (anonymous):

plz help ^-^

OpenStudy (ghostgate):

I believe you are correct, because if 1/8 in. was from the first week and 3/8 was from the second week then that means the plant grew 4/8 in. which divided by 2 = 2/4, divided by 2 again = 1/2 or 0.5.

OpenStudy (anonymous):

Thanks ^-^

OpenStudy (ghostgate):

You're welcome! I'm glad I could help!